OpenStudy (anonymous):

how to solve: 3/2 divided by 1/4=n. N is between?

OpenStudy (anonymous):

when you divide a fraction you can multiply it by flipping the fraction 3/2 * 4/1 = 12/2 = 6!!!!!

OpenStudy (anonymous):

Do you flip both fractions? or just the one?

OpenStudy (anonymous):

uno

OpenStudy (anonymous):

but it has to be the one at the end!!

OpenStudy (anonymous):

not the one in the front

OpenStudy (anonymous):

so you don't flip the first fraction, only the last?

OpenStudy (anonymous):

last

OpenStudy (anonymous):

perfect, thank you!

OpenStudy (anonymous):

yw